Circular No. 6360

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

COMET C/1996 B2 (HYAKUTAKE) W. M. Harris, F. Scherb, J. W. Percival and K. Nordsieck, University

of Wisconsin at Madison; and B. E. A. Mueller, D. Harmer and A. Macdonald, National Optical Astronomy Observatories; report the detection of a possible secondary condensation in the comet’s central region on a 70-s exposure on Mar. 25.47 UT using the imaging camera on the 3.5-m WIYN telescope at Kitt Peak: “The 7′ × 7′ field was obtained using a filter centered at 392 ± 10 nm under conditions of 0′′.6 seeing. To enhance detail the image was smoothed with a 4′′.0 (20 pixel) median filter and divided into the original frame. The divided result shows evidence for a separation of the nuclear region into two features. Both structures are located on the orbital path of the comet with a central separation of 0′′.66 or 51.2 km. The sunward of the two is ∼ 20-percent brighter than the other and displaced by ∼ 0′′.2 to the north. Because the separation of the structures is comparable to the seeing limit, they are not well enough resolved to determine if they are independent nuclei, although their 51-km separation argues against a single object. Both objects are 3–5 times brighter than the intensity of emission from the tail or any other jet structure observed in the coma.”

With respect to the observations above and those reported by J. Lecacheux et al. (IAUC 6354) and Tozzi et al. (IAUC 6357), as well as related information in the World-Wide Web (notably by H. A. Weaver regarding Hubble Space Telescope observations on Mar. 26.1 UT), Z. Sekanina, Jet Propulsion Laboratory, e-mails: “A tentative solution based on the available positions of what is apparently the same blob indicates that a fragment separated from the comet’s nucleus on Mar. 21.0±0.3 UT at a relative nongravitational deceleration of 280 ± 30 units of 10^{−5} solar attraction, typical for minor cometary companions whose lifetimes at 1 AU from the sun are usually 10–20 days. If it continues to survive, the piece should be located at p.a. 50^{o} from the nucleus, separated by 34′′ from it on Mar. 30.0 and by 45′′ on Apr. 5.0. This prediction refers to the tip of the feature nearest the nucleus. As the feature becomes ever more elongated, observers are likely to report average separations from the nucleus that will be increasingly greater than this prediction, by up to several arcseconds or more.”

Several readers, notably A. W. Harris, Jet Propulsion Laboratory, have pointed out that, if the tail of a comet corresponds to the extended radius vector (as is particularly likely for a gas tail), it is physically impossible for its angular extent to exceed the phase angle. They therefore question some of the reported observations of tail length, particularly on IAUC 6355. The phase angle of comet C/1996 B2 reaches a maximum of β = 111^{o} (IAUC 6359) around Apr. 6.

1996 March 28  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Brian G. Marsden